# Connections between Combinations Without Specified Separations and Strongly Restricted Permutations, Compositions, and Bit Strings

### v3

@@ 3.1 Connection with the class $\mathcal{Q}=\{1,q\}$
There is a bijection between the $k$-subsets of $\mathbb{N}_{n}$ such that no two elements of the subset differ by an element of $\mathcal{Q}=\{1,q\}$, where $q=2,3,\ldots$, and the permutations $\pi$ of $\mathbb{N}_{n+1}$ that have $k$ pairs of consecutive items that have between exchanged such that for all $i=1,\ldots,n+2-q$ the condition $\max_{j=0,\ldots,q-1}\pi(i+j)-\min_{j=0,\ldots,q-1}\pi(i+j)\leq q$ is met.
@@
The permutation $\pi$ of $\mathbb{N}_{n+1}$ corresponding to an allowed subset $\mathcal{S}$ of $\mathbb{N}_{n}$ where $\mathcal{Q}=\{1,q\}$ is formed as follows. If $i\in\mathcal{S}$ (where $i=1,\ldots,n$) then $\pi(i)=i+1$ and $\pi(i+1)=i$ (i.e., a pair of two consecutive items has been exchanged) whereas if neither $i$ nor $i-1$ are in $\mathcal{S}$ (for $i=2,\ldots,n+1$) then $\pi(i)=i$. As $\mathcal{Q}$ contains 1, once a pair of items has been exchanged, neither item will be moved again. Given that $\pi(i)$ can be no more than 1 away from $i$, the only way that the condition could be violated would be if $\pi(i+1)=i$ (and hence $\pi(i)=i+1$) and $\pi(i+q)=i+q+1$; this could only occur if both $i$ and $i+q$ were in $\mathcal{S}$ which is impossible since $q\in\mathcal{Q}$. The process of forming a permutation from a subset is clearly reversible and so the bijection is established. ∎
@@
Theorem 8 has connections to the following sequences in the OEIS [20]: A000930 ($S^{\{1,2\}}_{n}$) and A102547 ($S^{\{1,2\}}_{n,k}$), its associated triangle; A130137 ($S^{\{1,3\}}_{n}$), which we mention again in the section on bit strings; A263710 ($S^{\{1,4\}}_{n}$), which was the sequence that led to the discovery of the bijection and we will now prove the conjectured generating function for that entry; along with A374737 ($S^{\{1,5\}}_{n}$) and A385870 ($S^{\{1,6\}}_{n}$), which were recently added to the OEIS by the author.
The case $\mathcal{Q}=\{1,4\}$ falls into the class covered by the final theorem in [2], which gives recursion relations for $B^{\mathcal{Q}}_{n}$ and $B^{\mathcal{Q}}_{n,k}$. However, in order to avoid introducing the notation that theorem requires at this stage, we instead obtain the generating function using Theorem 5 which is also straightforward and illustrative to do.
Figure 3: Digraph for restricted-overlap tiling with squares ($S$) and $(2,2,1)$-combs ($C$), which corresponds to the $\mathcal{Q}=\{1,4\}$ case. In this and subsequent figures showing digraphs, each number in square brackets gives the contribution to the length of the metatile resulting from the addition of the tile. This subscript is omitted if there is no contribution, as is the case for any square placed in a gap of a comb.
@@
The $\mathcal{Q}$-comb is a $(2,2,1)$-comb. The metatile-generating digraph for tiling with squares and such combs (Fig. 3) has two inner cycles ($SC_{[3]}$ and $C_{[2]}C_{[3]}$) which have lengths $L_{1}=3$ and $L_{2}=5$. We choose $10^{2}$ 1 as the common node. The common circuits ($C_{[5]}S^{2}$ and $C_{[5]}C_{[2]}S$) have lengths $l_{\rm c1}=5$ and $l_{\rm c2}=7$. There is a single outer cycle ($S$) which has length $l_{\rm o1}=1$. Then from Theorem 5 and Corollary 7,
@@
The metatile-generating digraph corresponding to the next case in the $\{1,q\}$ series (Fig. 5) does not fall into any of the classes of digraph that we have derived a generating function for so far. However, as we will now show, the systematic application of the digraph generating function rules (Lemmas 2, 3, and 4) can still be used to yield the generating function in this instance. To assist in breaking the derivation into more manageable chunks, we introduce the following notation. We let $\operatorname{\mathit{G}_{X\to Y}^{\{Z\}}}$ denote the generating function corresponding to all possible walks from node $X$ to node $Y$ that do not include node $X$ a second time and never encounter node $Z$. We let $\operatorname{\mathit{G}_{X}^{\{Z\}}}$ denote the generating function corresponding to all possible walks from node $X$ to itself without visiting itself or node $Z$ in between.
@@
The generating function $G(x,y)$ corresponding to paths starting at the 0 node of the digraph in Figure 5 satisfies
@@
results from a loop within a loop: the $SC_{[4]}$ loop can be executed on arriving at the $10^{3}$ 1 node while traversing the $SC_{[4]}C_{[3]}$ loop. Combining the above results to give $G(x,y)$ and applying Corollary 7 to $G(x,1)$ yields (6) after simplifying. ∎
@@ 3.2 Connection with the class $\mathcal{Q}=\{m,2m,\ldots,jm\}$
The proof we give of the second bijection (Theorem 13) requires the following bijection established in [8] between strongly restricted permutations and tiling with $(\frac{1}{2},g)$-fences (which are just $(\frac{1}{2},g,\frac{1}{2})$-combs) where $g$ is always a nonnegative integer. As the teeth (originally called posts in the context of fences) are not of integer width, it is convenient to regard each cell as being divided into two slots into which a tooth can fit [9].
@@
There is a bijection between (i) the permutations of $\mathbb{N}_{n}$ satisfying $\pi(i)-i\in\mathcal{D}$ for each $i\in\mathbb{N}_{n}$ and (ii) the tilings of an $n$-board using $(\frac{1}{2},d_{j})$-fences with the left tooth always placed in a left slot and $d_{j}$ equal to the non-negative elements of $\mathcal{D}$ and $(\frac{1}{2},-d_{k}-1)$-fences with their gaps aligned with cell boundaries and $d_{k}$ equal to the negative elements of $\mathcal{D}$. If the left slot of cell $i$ of the $n$-board is occupied by the left (right) tooth of a $(\frac{1}{2},g)$-fence then $\pi(i)=i+g$ ($\pi(i)=i-g-1$). For example, the tiling in Fig. 6 corresponds to the permutation
@@
We refer to a fence corresponding to an excedance (i.e., a position $i$ such that $\pi(i)>i$) as an up fence and a fence giving $\pi(i)-i<0$ as a down fence. The case of a fixed point ($\pi(i)=i$) corresponds to a gapless fence which is just an ordinary square tile (that must be aligned with a cell).
The following bijection is an extension of that noted but not proved by Baltić [6]. The proof we give requires the notion of a comb with $t$ teeth all of length $w$ separated by gaps of length $g$ which we refer to as a $(w,g;t)$-comb [5, 4].
@@
Corollary 11 in [1], for $l\geq 0$, $m\geq 1$, $t\geq 2$, and $r=0,\ldots,m-1$, gives us
@@
As a consequence of the bijection given in Theorem 13, $P^{\{-m,0,jm\}}_{n+jm}=S^{\{m,2m,\ldots,jm\}}_{n}$ and so $P^{\{-m,0,jm\}}_{n}=S^{\{m,2m,\ldots,jm\}}_{n-jm}$. Replacing $n$ by $im+r$ in this and using (9) with $t$ replaced by $j+1$ and $l$ replaced by $i-t+1$ gives (7). Similarly, Corollary 10 in [1] gives
@@ 4 Connections with compositions
A composition of a positive integer $n$ is a way of expressing $n$ as a sum of positive integers (referred to as parts) where the order of the parts is significant; thus, for example, 1+2 and 2+1 are different compositions of 3. For convenience, it is customary to take the number of compositions of 0 (any negative integer) as being 1 (0). It is well known that the number of compositions of $n$ is the same as the number of tilings of an $n$-board using $m$-ominoes for arbitrary nonnegative integers $m$ and hence that the number of compositions of $n$ into distinct parts $m_{1},m_{2},\ldots,m_{i},\ldots$ is given by $B_{n}=\delta_{n,0}+\sum_{i=1}B_{n-m_{i}}$, $B_{n<0}=0$ for all integers $n$. It follows that any tiling in which all possible metatiles have distinct lengths corresponds to a composition into parts equal to the lengths of the metatiles. We first examine two classes of $\mathcal{Q}$ where there are a finite number of possible metatiles when restricted-overlap tiling with squares and $\mathcal{Q}$-combs. Recall that the number of possible metatiles is finite iff $2r\geq q$, where $r$ is the length of the rightmost tooth of the $\mathcal{Q}$-comb (Lemma 2 in [2]). It is sometimes more convenient to express this condition as $r\geq s-1$, where $s=q+1-r$ is the sum of the lengths of the teeth and gaps up to the rightmost tooth.
@@
Suppose $q>0$ is the largest element of $\mathcal{Q}$. Then $S^{\mathcal{Q}}_{n}$ is the number of compositions of $n+q$ into parts which are drawn from a finite set of distinct parts iff every metatile, when restricted-overlap tiling with squares and $\mathcal{Q}$-combs, contains no more than two combs. Then the parts are 1, $q+1$, and $q+1+p_{i}$, where $p_{i}$ is the cell number of the $i$-th empty cell in the $\mathcal{Q}$-comb.
@@
If there are a finite number of parts, there are a finite number of metatiles which means that, in all cases, the start of the final comb in a metatile containing more than one comb must lie within the first comb (see the proof of Lemma 2 in [2]). If there were a third comb present, the cells it occupies without overlapping (of which there must be at least one, i.e., its cell 0) could be replaced by squares and give rise to a metatile of the same length. The parts would then not be distinct. The square and filled comb metatiles are of length 1 and $q+1$, respectively. The remaining metatiles contain two combs. If the start of the second comb is in cell $p_{i}$ of the first comb, the resulting metatile (completed by filling any remaining empty cells with squares) is easily seen to have a length of $q+1+p_{i}$. ∎
In the proofs of the following two theorems we employ combs that, in general, have a periodic pattern of teeth and gaps before the rightmost tooth is reached. Extending the notation used for combs with teeth all the same length and gaps all the same length [4, 5], we call an $(l,g,l,g,\ldots,r)$-comb with $t$ teeth an $(l,g,r;t)$-comb and we refer to an $(l,g,m,h,l,g,m,h,\ldots,r)$-comb with $t$ teeth as an $(l,g,m,h,r;t)$-comb. Note that an $r$-omino (an $(l,g,r)$-comb) can be regarded as an $(l,g,r;t)$- or $(l,g,m,h,r;t)$-comb with $t=1$ ($t=2$).
@@
Here the $\mathcal{Q}$-comb is an $(l,g,m,h,r;t)$-comb. The $t=1,2$ cases are covered by the proof of Theorem 16. First note that if $t$ is odd, the length of the comb $q+1=\frac{1}{2}(t-1)(l+g+m+h)+r$ and it is $\frac{1}{2}t(l+g)+(\frac{1}{2}t-1)(m+h)+r$ if $t$ is even from which the respective conditions on $r$ (so that there are a finite number of metatiles) follow. There are two mutually exclusive sets of conditions that ensure that the metatiles contain no more than two combs. Condition (i) that $l\geq g+m+h$ (see Fig. 7(b)) arises as a result of the observation that if $l$ were one less than $g+m+h$ and a second comb were placed starting at the first gap of the first comb, the start a third comb could be placed starting at the final cell of the second gap of the first comb. The first part of condition (ii), namely that $l\geq g,h$ is for the same reasons as the $l\geq g$ condition in Theorem 16. We must also have $l+g+m\geq g+m+h$ (see Fig. 7(c)) which also implies that $l\geq h$. Were $l+g+m$ one less than $g+m+h$, it can be seen that after placing the second comb at the start of the first gap, a third comb could be placed at the end of the second gap of the first comb. At the same time, we must ensure that $m$ is not too small; otherwise the start of a third comb could be inserted at the start of the second gap of the first comb. Thus $l-1+g\leq m$ (see Fig. 7(d)). The result then follows from Lemma 15. ∎
We now turn to compositions where the parts are drawn from an infinite set. In the context of tiling this means that there is an infinite number of possible metatiles and so the metatile-generating digraph has an inner cycle.

### v4

@@ 3.1 Connection with the class $\mathcal{Q}=\{1,q\}$
There is a bijection between the $k$-subsets of $\mathbb{N}_{n}$ such that no two elements of the subset differ by an element of $\mathcal{Q}=\{1,q\}$, where $q=2,3,\ldots$, and the permutations $\pi$ of $\mathbb{N}_{n+1}$ that have $k$ pairs of consecutive items that have been exchanged such that for all $i=1,\ldots,n+2-q$ the condition $\max_{j=0,\ldots,q-1}\pi(i+j)-\min_{j=0,\ldots,q-1}\pi(i+j)\leq q$ is met.
@@
The permutation $\pi$ of $\mathbb{N}_{n+1}$ corresponding to an allowed subset $\mathcal{S}$ of $\mathbb{N}_{n}$ where $\mathcal{Q}=\{1,q\}$ is formed as follows. If $i\in\mathcal{S}$ (where $i=1,\ldots,n$) then $\pi(i)=i+1$ and $\pi(i+1)=i$ (i.e., a pair of two consecutive items has been exchanged) whereas if neither $i$ nor $i-1$ are in $\mathcal{S}$ (for $i=2,\ldots,n+1$) then $\pi(i)=i$. As $\mathcal{Q}$ contains 1, once a pair of items has been exchanged, neither item will be moved again. Given that $\pi(i)$ can be no more than 1 away from $i$, the only way that the condition could be violated would be if $\pi(i+1)=i$ (and hence $\pi(i)=i+1$) and $\pi(i+q)=i+q+1$; this could only occur if both $i$ and $i+q$ were in $\mathcal{S}$. This is impossible since $q\in\mathcal{Q}$. The process of forming a permutation from a subset is clearly reversible and so the bijection is established. ∎
@@
Theorem 8 has connections to the following sequences in the OEIS [21]: A000930 ($S^{\{1,2\}}_{n}$) and A102547 ($S^{\{1,2\}}_{n,k}$), its associated triangle; A130137 ($S^{\{1,3\}}_{n}$), which we mention again in the section on bit strings; A263710 ($S^{\{1,4\}}_{n}$), which was the sequence that led to the discovery of the bijection and we now prove the conjectured generating function for that entry; along with A374737 ($S^{\{1,5\}}_{n}$) and A385870 ($S^{\{1,6\}}_{n}$), which the author recently added to the OEIS.
The case $\mathcal{Q}=\{1,4\}$ falls into the class covered by Theorem 8 in All24, which gives recursion relations for $B^{\mathcal{Q}}_{n}$ and $B^{\mathcal{Q}}_{n,k}$. However, in order to avoid introducing the notation used in that theorem at this stage, we instead obtain the generating function using Theorem 5, which is also straightforward and illustrative to do.
Figure 3: Digraph for restricted-overlap tiling with squares ($S$) and $(2,2,1)$-combs ($C$), which corresponds to the $\mathcal{Q}=\{1,4\}$ case. In this and subsequent figures showing digraphs, each number in square brackets gives the contribution to the length of the metatile resulting from the addition of the tile. The subscript is omitted if there is no contribution, as is the case for any square placed in a gap of a comb.
@@
The $\mathcal{Q}$-comb is a $(2,2,1)$-comb. The metatile-generating digraph for tiling with squares and such combs (Fig. 3) has two inner cycles ($SC_{[3]}$ and $C_{[2]}C_{[3]}$). These have lengths $L_{1}=3$ and $L_{2}=5$. We choose $10^{2}$ 1 as the common node. The common circuits ($C_{[5]}S^{2}$ and $C_{[5]}C_{[2]}S$) have lengths $l_{\rm c1}=5$ and $l_{\rm c2}=7$. There is a single outer cycle ($S$). This has length $l_{\rm o1}=1$. Then from Theorem 5 and Corollary 7 we get
@@
The metatile-generating digraph corresponding to the next case in the $\{1,q\}$ series (Fig. 5) does not fall into any of the classes of digraph that we have derived a generating function for so far. However, as we now show, the systematic application of the digraph generating function rules (Lemmas 2, 3, and 4) can still be used to yield the generating function in this instance. To assist in breaking the derivation into more manageable chunks, we introduce the following notation. We let $\operatorname{\mathit{G}_{X\to Y}^{\{Z\}}}$ denote the generating function corresponding to all possible walks from node $X$ to node $Y$ that do not include node $X$ a second time and never encounter node $Z$. We let $\operatorname{\mathit{G}_{X}^{\{Z\}}}$ denote the generating function corresponding to all possible walks from node $X$ to itself without visiting itself or node $Z$ in between.
@@
The generating function $G(x,y)$ corresponding to walks starting at the 0 node of the digraph in Figure 5 satisfies
@@
results from a loop within a loop: the $SC_{[4]}$ loop can be executed on arriving at the $10^{3}$ 1 node while traversing the $SC_{[4]}C_{[3]}$ loop. Combining the above results to give $G(x,y)$ and applying Corollary 7 yields (6) after simplifying. ∎
@@ 3.2 Connection with the class $\mathcal{Q}=\{m,2m,\ldots,jm\}$
The proof we give of the second bijection (Theorem 13) requires the following bijection (established in previous work [9]) between strongly restricted permutations and tiling with $(\frac{1}{2},g)$-fences (which are just $(\frac{1}{2},g,\frac{1}{2})$-combs), where $g$ is always a nonnegative integer. As the teeth (originally called posts in the context of fences) are not of integer width, it is convenient to regard each cell as being divided into two slots into which a tooth can fit [10].
@@
There is a bijection between (i) the permutations of $\mathbb{N}_{n}$ satisfying $\pi(i)-i\in\mathcal{D}$ for each $i\in\mathbb{N}_{n}$ and (ii) the tilings of an $n$-board using $(\frac{1}{2},d_{j})$-fences with the left tooth always placed in a left slot and $d_{j}$ equal to the non-negative elements of $\mathcal{D}$ and $(\frac{1}{2},-d_{k}-1)$-fences with their gaps aligned with cell boundaries and $d_{k}$ equal to the negative elements of $\mathcal{D}$. If the left slot of cell $i$ of the $n$-board is occupied by the left (right) tooth of a $(\frac{1}{2},g)$-fence then $\pi(i)=i+g$ ($\pi(i)=i-g-1$).
For example, the tiling in Fig. 6 corresponds to the permutation
@@
We refer to a fence corresponding to an excedance (i.e., a position $i$ such that $\pi(i)>i$) as an up fence and a fence giving $\pi(i)-i<0$ as a down fence. The case of a fixed point ($\pi(i)=i$) corresponds to a gapless fence. This is just an ordinary square tile (that must be aligned with a cell).
The following bijection is an extension of that noted but not proved by Baltić [7]. The proof we give requires the notion of a comb with $t$ teeth all of length $w$ separated by gaps of length $g$, which we refer to as a $(w,g;t)$-comb [4, 5].
@@
Using a previously obtained result [1, Corollary 11], for $l\geq 0$, $m\geq 1$, $t\geq 2$, and $r=0,\ldots,m-1$, we have
@@
As a consequence of the bijection given in Theorem 13, $P^{\{-m,0,jm\}}_{n+jm}=S^{\{m,2m,\ldots,jm\}}_{n}$ and so $P^{\{-m,0,jm\}}_{n}=S^{\{m,2m,\ldots,jm\}}_{n-jm}$. Replacing $n$ by $im+r$ in this and using (9) with $t$ replaced by $j+1$ and $l$ replaced by $i-t+1$ gives (7). Similarly, using another result from previous work [1, Corollary 10], we obtain
@@ 4 Connections with compositions
A composition of a positive integer $n$ is a way of expressing $n$ as a sum of positive integers (referred to as parts) where the order of the parts is significant; thus, for example, 1+2 and 2+1 are different compositions of 3. For convenience, it is customary to take the number of compositions of 0 (any negative integer) as being 1 (0). It is well known that the number of compositions of $n$ is the same as the number of tilings of an $n$-board using $m$-ominoes for arbitrary nonnegative integers $m$ and hence that the number of compositions of $n$ into distinct parts $m_{1},m_{2},\ldots,m_{i},\ldots$ is given by $B_{n}=\delta_{n,0}+\sum_{i=1}B_{n-m_{i}}$, $B_{n<0}=0$ for all integers $n$. It follows that any tiling in which all possible metatiles have distinct lengths corresponds to a composition into parts equal to the lengths of the metatiles.
4.1 Compositions with parts drawn from finite sets
We first examine two classes of $\mathcal{Q}$ where there are a finite number of possible metatiles (of distinct lengths) when restricted-overlap tiling with squares and $\mathcal{Q}$-combs. Recall that the number of possible metatiles is finite iff $2r\geq q$, where $r$ is the length of the rightmost tooth of the $\mathcal{Q}$-comb [2, Lemma 2]. It is sometimes more convenient to express this condition as $r\geq s-1$, where $s=q+1-r$ is the sum of the lengths of the teeth and gaps up to the rightmost tooth.
@@ 4.1 Compositions with parts drawn from finite sets
Suppose $q>0$ is the largest element of $\mathcal{Q}$. Then $S^{\mathcal{Q}}_{n}$ is the number of compositions of $n+q$ into parts which are drawn from a finite set of distinct parts iff every metatile, when restricted-overlap tiling with squares and $\mathcal{Q}$-combs, contains no more than two combs. Then the parts are 1, $q+1$, and $q+1+p_{i}$ (for $i=1,\ldots,\left\lvert\mathbb{N}_{q}\setminus\mathcal{Q}\right\rvert$), where $p_{i}$ is the cell number of the $i$-th empty cell in the $\mathcal{Q}$-comb.
@@
If there are a finite number of parts, there are a finite number of metatiles. This means that, in all cases, the start of the final comb in a metatile containing more than one comb must lie within the first comb (see the proof of Lemma 2 in All24). If there were a third comb present, the cells it occupies without overlapping (of which there must be at least one, i.e., its cell 0) could be replaced by squares and give rise to a metatile of the same length. The parts would then not be distinct. The square and filled comb metatiles are of length 1 and $q+1$, respectively. The remaining metatiles contain two combs. If the start of the second comb is in cell $p_{i}$ of the first comb, the resulting metatile (completed by filling any remaining empty cells with squares) is easily seen to have a length of $q+1+p_{i}$. ∎
In the proofs of the following two theorems we employ combs that, in general, have a periodic pattern of teeth and gaps before the rightmost tooth is reached. Extending the notation used for combs with teeth all the same length and gaps all the same length [4, 5], we call an $(l,g,l,g,\ldots,r)$-comb with $t$ teeth an $(l,g,r;t)$-comb and we refer to an $(l,g,m,h,l,g,m,h,\ldots,r)$-comb with $t$ teeth as an $(l,g,m,h,r;t)$-comb. Note that an $r$-omino (an $(l,g,r)$-comb) can be regarded as an $(l,g,r;t)$- or $(l,g,m,h,r;t)$-comb with $t=1$ ($t=2$). Recall that the inclusion of 0 in $\mathcal{Q}$ is merely for convenience in the statement of the theorems and is to be removed when giving particular instances of $\mathcal{Q}$ that the theorems apply to.
@@
Here the $\mathcal{Q}$-comb is an $(l,g,m,h,r;t)$-comb. The $t=1,2$ cases are covered by the proof of Theorem 16. First note that if $t$ is odd, the length of the comb $q+1=\frac{1}{2}(t-1)(l+g+m+h)+r$ and it is $\frac{1}{2}t(l+g)+(\frac{1}{2}t-1)(m+h)+r$ if $t$ is even from which the respective conditions on $r$ (so that there are a finite number of metatiles) follow. There are two mutually exclusive sets of conditions that ensure that the metatiles contain no more than two combs. Condition (i) that $l\geq g+m+h$ (see Fig. 7(b)) arises as a result of the observation that if $l$ were one less than $g+m+h$ and a second comb were placed starting at the first gap of the first comb, the start a third comb could be placed starting at the final cell of the second gap of the first comb. The first part of condition (ii), namely that $l\geq g,h$ is for the same reasons as the $l\geq g$ condition in Theorem 16. We must also have $l+g+m\geq g+m+h$ (see Fig. 7(c)). This also implies that $l\geq h$. Were $l+g+m$ one less than $g+m+h$, it can be seen that after placing the second comb at the start of the first gap, a third comb could be placed at the end of the second gap of the first comb. At the same time, we must ensure that $m$ is not too small; otherwise the start of a third comb could be inserted at the start of the second gap of the first comb. Thus $l-1+g\leq m$ (see Fig. 7(d)). The result then follows from Lemma 15. ∎
4.2 Compositions with parts drawn from infinite sets
We now turn to compositions where the parts are drawn from infinite sets. In the context of tiling this means that there is an infinite number of possible metatiles and so the metatile-generating digraph has an inner cycle. We start by showing that the metatile lengths can only be all different if there is just one inner cycle.